Chapter 5: A Survey of Probability ConceptsISBN : 0072868244 Author : Douglas A. Lind, William G. Marchal, Samuel A. Wathen
copyright © 2004 The McGraw-Hill Companies
exercise 42;
A telephone number consists of seven digits, the first three representing the exchange. How many different telephone numbers are possible within the 537 exchange.

Found 2 solutions by ManMan, mathslover:
Answer by ManMan(1)

How many different telephone numbers are possible within the 537 exchange.
537 _ _ _ _
The blanks can be the numbers: 0 1 2 3 4 5 6 7 8 OR 9.
There are 10 choices in total; Multiply the numbers to get your answer:
10*9*8*7*6*5*4*3*2*1
10*9=90
90*8=720
720*7=5040
5040*6=30240
30240*5=151200
151200*4=604800
604800*3=1814400
1814400*2=3628800
3628800*1=3628800
The final answer is: 3,628,800

Answer by mathslover(132)

we have the number as 537_ _ _ _
Each of the blanks can be filled in 10 ways ( either of the numbers from 0-9)
So total numbers of ways = 10 * 10 * 10 * 10
=10,000
Alternately, the numbers will be from 537 0000 to 537 9999
clearly there are 10000 different such numbers